OpenStudy (anonymous):

what did I do wrong? solve for x -10e^(2-2x)-6=-66

OpenStudy (anonymous):

\[-10e ^{(2-2x)}-6=-66\]\[-10e ^{(2-2x)}=-60\]\[e ^{(2-2x)}=6\]\[\ln e ^{(2-2x)}=\ln 6\]\[2-2x=\ln 6\]\[-2x=(\ln 6)-2\]\[x=\frac{ ((\ln 6)-2) }{ -2 }\]